Suppose that the functions p and p(x) = x² + 6 g(x)=Vx+1 Find the following. (P 9 g)(3) = [] (40p) (3) = 0 Х
Answer #1

1.

(p.q)(x) = p(q(x)) = p((x+1)) = ((x+1)2 + 6 = x + 7

(p.q)(3) = 3 + 7= 10

2.

(q.p)(x) =  q(p(x)) = q (x2+6) = ((x2+6)+1) = (x2+7)

(q.p)(3) = (32+7) = (16) = 4